OpenStudy (ikou):

What is the simplified form of the quantity 9 x squared minus 25 over the quantity 3 x plus 5?

OpenStudy (ikou):

https://gyazo.com/660e4524c7935264f2b85dc266e776c8 what the equation looks like

OpenStudy (anonymous):

The numerator is a conjugate: \[9x ^{2}-25 = (3x + 5)(3x - 5)\] Now look at your denominator. What happens when a term is divided by its identical self? @Ikou

OpenStudy (ikou):

would it equal zero?

OpenStudy (anonymous):

Not quite. What is any number divided by the same number? For example: \[\frac{ 2 }{ 2 } = ~?\]

OpenStudy (ikou):

1?

OpenStudy (anonymous):

Precisely. Likewise: \[\frac{ 3x + 5 }{ 3x + 5 } = ~?\]

OpenStudy (ikou):

it would equal one

OpenStudy (anonymous):

Exactly. Nice job. What happens when you multiply by one? \[1 \times (3x - 5) = ~?\]

OpenStudy (ikou):

It's still gonna be 3x-5

OpenStudy (anonymous):

You hit the jackpot @Ikou. That's the simplified form of your initial expression. ^_^

OpenStudy (ikou):

Oh oki, so which one would be the answer exactly? https://gyazo.com/c1fd2eb916e5d6238ca0ed15f9a3c09b

OpenStudy (anonymous):

Okay, the 'restriction' those are referring to is the denominator. Specifically, when the denominator becomes zero, the entire function is considered undefined. So, to find this value, we take our denominator (3x + 5) from the original equation and equate that to zero. Like this: \[3x + 5 = 0\] Solve for x. \[x = -\frac{ 5 }{ 3 }\] Which means that for our function: \[x \neq -\frac{ 5 }{ 3 }\] This is our restriction.

OpenStudy (ikou):

Did I cross out the answer?

OpenStudy (ikou):

I crossed out all the '+'

OpenStudy (anonymous):

Nope, you left it open. It's C. You need the 3x-5 we got earlier and the restriction x cannot equal -5/3.

OpenStudy (ikou):

Oh okii, well thank you for all your help!

OpenStudy (anonymous):

You're welcome